Observing FLAO_SX+Luci1

Staff:

Remote control room: Juan Carlos Guerra, Doug Miller

From Home: Dave Thompson.

Telescope Operator: David Gonzalez.

Summary

  • Hand over from ARGOS about 12am
  • Observe in seeing limited with Luci1 in close loop.
  • Determination of the Image quality of Luci in; active optics,TTM and ACE mode.
  • Target to observe M103.1:33:14,+60:41:12
  • Script: /home/observer/LBTO/DThompson/AO_TTM/M103_TTM.img
  • Seeing good <0.9" and stable.
  • Clear sky.
  • Luci Images;Luci.20141001.

WFS_SX

  • Startup .Ok
  • Hardware.Ok

AdsecSX

We had a few RIP due to higher force in actuators during the close loop.

Details


Test ACE preset to a bright star, the star is in the catalogue; AO799 with magnitude 9.
Saved telemetry data of the close loop:
20141001_074117, 20141001_074208,_074311

The seeing is stable <0.8" . The differential focus between the pyramid and the the Active optics is found to be 53.5mm in the pyramid z stage. The new value is updated in the file acceptance in the wfs side. For the record the previous value was 44.5mm.

Preset to Messier 103

The preset in done in TTM which load only 5 correcting modes. Telemetry data were saved;20141001_083333,_08350

The mirror rip 20140110 09:12:09.9457 "Alert: Error: COILS DISABLED". The loop is open, The Adsec remains in AORunning status,WFS status is AOSet, and it was stuck executing Recover Failure, also the elevation telemetry value was not available. The system took long time to recover, the adsecArbitrator was stop/start and it didn't clean the Adsec Status after two minutes the system complete the recover failure and the shell can be seeing in rest position.

Sequence:

TTM

0011-0015 8x4.0s BrG G*5 @ 11.77m TTM(5 modes, 400Hz)

FWHM ~0.36" - Dave comment: Doesn't really look any better than the seeing-ltd data

The loop is stable. The optical loop telemetry is saved;20141001_083333,_083508

ACE

0016-0020 8x4.0s BrG G*5 @ 11.77m ACE (110 modes, 400Hz, bin2).

loop stable. The optical loop telemetry is saved;_093004,_093049

TTM

0021-0025 8x4.0s BrG G*5 @ 11.77m TTM ( 5 modes, 400Hz, bin4).

Loop stable. The optical loop telemetry is saved;_094228,_094305,_094335,_094841

Seeing limited (off axis)


0026-0030 8x4.0s BrG G*5 @ 11.77m Seeing-Limited (off axis GS 32)

ACE (bright)

0031-0033 8x4.0s BrG G*0 @ 7.51m ACE (400 modes, 990Hz, bin1 - "the best")

The loop opened in third exposure. Forces too high on 3 actuators.
Script failed out because 4th offset failed.

Loop is stable with some spikes of -0.56N in force. The optical loop telemetry is saved; _100412,_100511

ACE (faint)

0034-0038 8x4.0s BrG G*14 @ 15.70m ACE (36 modes, 100Hz, bin4)

The magnitude measured in the FLAO is 16.6. The first try to close the loop failed and the second attemp was succesful.

The loop remains stable with about 20photons/subap.

Optical loop telemetry saved: _102208,_102228,_102413,_103129

From Dave log:

> 07:41 0.6" on AGw guider (23:37 +26:11)
>
> Target for TTM test is Messier 103 at 01:33 +60
> See script: /home/observer/LBTO/DThompson/AO_TTM/M103_TTM.img
>
> First execution of script...error
> Pause(FLAO) - AOARB: Illegal command for state ReadyToAcquire (Pause)
>
> 0001 0005 8x4.0s BrG Seeing-limited mode. 0.48" on guider 0.4" on LUCI
> 0006 0010 8x4.0s BrG G*5 @ 11.77m TTM(5 modes, 400Hz)
> Images 2-5 of sequence aberrated...apparently the system reverted to the
> initial/default focus value. JC correcting the config file.
>
> 0011 0015 8x4.0s BrG G*5 @ 11.77m TTM(5 modes, 400Hz)
> FWHM ~0.36" - Doesn't really look any better than the seeing-ltd data
>
> Shell ripped when setting up on next star...
>
> 0016 0020 8x4.0s BrG G*5 @ 11.77m ACE (110 modes, 400Hz, bin2)
> 0021 0025 8x4.0s BrG G*5 @ 11.77m TTM ( 5 modes, 400Hz, bin4)
> 0026 0030 8x4.0s BrG G*5 @ 11.77m Seeing-Limited (off axis GS 32)
> 0031 0033 8x4.0s BrG G*0 @ 7.51m ACE (400 modes, 990Hz, bin1 - "the best")
> Loops opened in third exposure. Forces too high on 3 actuators.
> Script failed out because 4th offset failed.
>
> ~0.55" on guider
> 0034 0038 8x4.0s BrG G*14 @ 15.70m ACE (36 modes, 100Hz, bin4)
> Loop would not close on first attempt.
> Star measured as 16.6 mag on pyramid

Comments


 
Topic revision: r1 - 01 Oct 2014, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback